OpenStudy (anj123):

What is the integral of [(tanx)^3(secx)^6]dx ?

Nnesha (nnesha):

\[\large\rm \int\limits_{ }^{ } (tanx)^3 (secx)^6 dx\]

Nnesha (nnesha):

option: break sec^6x nto sec^2x or option 2: rewrite tan^3x as tan^1x

Nnesha (nnesha):

\[\huge\rm \int\limits_{ }^{} tanx \cdot \tan^2x \cdot ( secx)^6 dx\] let u =secx

Nnesha (nnesha):

\[\large\rm u = \sec x , ~~du=secx ~tanx~ dx\]u=secx so we can replace (secx)^6 with u\[\large\rm \int\limits_{ }^{ } \tan x \cdot \tan^2x \cdot u^6 \frac{du}{secx~tanx}\] remember the identity 1+tan^2x=sec^2x solve for tan^2x we will get `tan^2x =sec^2x-1` plug in for tan^2x u^6 can be written as u times u^5 right ? \[\large\rm \int\limits_{ }^{ } \tan x \cdot \tan^2x \cdot u \cdot u^5 \frac{du}{secx~tanx}\] replace u =secx \[\large\rm \int\limits_{ }^{ } \tan x \cdot \color{Red}{\tan^2x} \cdot (secx)\cdot u^5 \frac{du}{secx~tanx}\] \[\large\rm \int\limits_{ }^{ } \tan x \cdot \color{Red}{(sec^2x-1)} \cdot (secx)\cdot u^5 \frac{du}{secx~tanx}\] \[\large\rm \int\limits_{ }^{ } \tan x\cdot (secx) \cdot \color{Red}{(sec^2x-1)} \cdot u^5 \frac{du}{secx~tanx}\] now we cancel out some *stuffs* make sense or no ??

Nnesha (nnesha):

can*

OpenStudy (anj123):

\[\frac{ (secx)^8 }{ 8 }-\frac{ (secx)^6 }{ 6 } +C\]

Nnesha (nnesha):

correct

OpenStudy (anj123):

Is that the answer?

OpenStudy (anj123):

okay. Thanks

Nnesha (nnesha):

yw